# Physical Processes Behind the Co-Evolution of Halos, Galaxies and Supermassive Black Holes in the IllustrisTNG Simulation

### v2

@@ 3 Evolutionary Phases and Transitions
Figure 2: The phases and transitions in the evolution histories of galaxies. Left panel shows the $M_{\rm BH}$-$M_{*}$ histories of galaxies. Grey curves represent the simulated histories of 20 individual galaxies randomly selected from our sample (see §2). Dots are the history of an example galaxy, colored blue (red) if it is star-forming (quenched) at a snapshot. Red dashed curve shows the analytical fitting of the history for this example. Four phases are defined according to the fitting function and are labeled at the top of the panel. Three time points, $t_{\rm 12}$, $t_{\rm 23}$ and $t_{\rm 34}\equiv t_{\rm q}$, are used to separate these phases (see §3 for the fitting method and the definition of phases and transitions). Black curve shows the median $M_{\rm BH}$-$M_{*}$ relation at $z=0$ computed using all galaxies in our sample, with the shaded area indicating the $16^{\rm th}$$84^{\rm th}$ percentiles. Short black line indicates a power law, $M_{\rm BH}\propto M_{*}^{5/3}$, to which the regulation of thermal-mode AGN feedback in Phase-3 tends to drive the galaxy (see §4.3). Right panel shows the cumulative distribution of $\tau_{i}$, the duration of Phase-$i$.
The similarity of different galaxies shown in Fig. 1 indicates that a universal function can be used to fit the evolution path of each galaxy in the $M_{\rm BH}$-$M_{*}$ plane. For each galaxy, we identify its quenching time, $t_{\rm q}$, defined as the first time when its ${\rm sSFR}(t)$, smoothed by a running Gaussian kernel with standard deviation equal to the dynamical time scale ($t_{\rm dyn,h}\equiv R_{\rm h}/V_{\rm h}$) of the host halo at each snapshot, is less than $10^{-11}\,\,{\rm yr}^{-1}$. Then, we fit all snapshots before $t_{\rm q}$ by the following function
@@
where $y=\log(M_{\rm BH}/\,{\rm M_{\odot}})$, $x=\log(M_{\rm*}/\,{\rm M_{\odot}})$, and $(x_{0},y_{0},k,a,b)$ are the five free parameters characterizing the evolutionary phases and transitions. The values of the free-parameters are obtained by minimizing the loss function, $\chi^{2}=\sum_{s}(y_{{\rm sim},s}-y_{s})^{2}$, computed using the simulated value ($y_{\rm sim}$) and the value predicted by the fitting function ($y$) at every snapshot, $s$. The optimization is performed using the Levenberg-Marquardt algorithm, which is designed for solving non-linear least squares problems. By design, the fitting function explicitly includes three transition points at which the behavior of the galaxy in the $M_{\rm BH}$-$M_{*}$ plane changes significantly. The first transition point is $x=x_{0}-b$ at which ${\rm d}y/{\rm d}x=1$. This marks the time when the growth of the SMBH begins to outpace the star formation. The second transition is at $x=x_{0}$, where the rapid growth in the SMBH mass ends. Given that TNG adopts a Bondi accretion upper-bound only by the Eddington limit, the growth of SMBH during $x_{0}-b<x<x_{0}$ can be very fast. This behavior can be captured by a value of $a$ close to zero. For $x\geqslant x_{0}$, the function $y(x)$ is designed to be linear, describing a power-law relation, $M_{\rm BH}\propto M_{*}^{k}$. The final transition is the ending point of the domain of the fitting function, which, by design, is the quenching point of the galaxy. The best-fit function for an example galaxy is shown by the red dashed curve in the left panel of Fig. 2, with three black hexagons marking the three transition points (see below for the notations), respectively. The fitting captures well all the nonlinear features of the simulated points in this example. For reference, in Appendix A, we quantify the goodness-of-fit for all galaxies in our sample. For most galaxies in our sample, the root-mean-square error is less than $0.1\,{\rm dex}$, and the maximum absolute error is less than $0.3\,{\rm dex}$, indicating that the fittings are reliable.
@@
  • Phase-4: Merger dominated growth. This phase is characterized by the cease of in-situ star formation and SMBH growth. Consequently, the trajectory of a galaxy in the $M_{\rm BH}$-$M_{*}$ plane is stuck at around a point near the scaling relation at $z=0$, with some occasional “jumps” due to mergers.
@@ 4 Processes Driving the Evolution
Figure 3: Evolution history of various properties at around the transition point $t_{12}$. Each property describing the history is averaged over our sample (see §2) and shown as a function of time, measured with respect to $t_{12}$ and scaled by the duration $\tau_{2}+\tau_{3}$. The shaded area around each curve shows the $16^{\rm th}$$84^{\rm th}$ percentiles. Each panel shows a set of related properties. Top left panel shows the specific growth rates of of $M_{\rm h}$ (blue), $M_{*}$ (orange) and $M_{\rm BH}$ (green). Top right panel shows the gas properties involved in the definition of Bondi accretion rate (see Eq. 3 and 4.1) in the vicinity of the central SMBH. Three bottom panels show the gas masses defined at different scales: bound to the subhalo (left), within $2R_{\rm*,1/2}$ (middle), and within $0.2R_{\rm*,1/2}$ (right). Total, cold ($\leqslant 10^{5}\,{\rm K}$), and hot ($>10^{5}\,{\rm K}$) gas masses are shown by black, blue and red curves, respectively, all scaled by total gas mass at $t_{12}$.
Figure 4: Relation of the estimated value and the simulated value of the Phase-1 duration ($\tau_{1}$). Grey curves, from inner to outer, are contours enclosing 10% – 95% of the sample. Dashed line is a power-law fitting (in linear scale) of the relation. The normalization and power index are both close to 1, indicating that the estimated value well approximates the simulated value. See §4.1 for the details.
@@ 4.1 Phase-1: Star-formation dominated growth
The growth of a galaxy in the first phase is featured by a small mass of the SMBH. In TNG, a seed of SMBH with a mass of $M_{\rm BH}=1.18\times 10^{6}\,{\rm M_{\odot}}$ is placed into a galaxy when the mass of its host halo exceeds $7.38\times 10^{10}\,{\rm M_{\odot}}$ (Weinberger et al., 2018). The subsequent growth of the SMBH is modeled by the Bondi-Hoyle accretion (Weinberger et al., 2017, thereafter Bondi accretion). As the accretion rate of the SMBH modeled in this way scales as $M_{\rm BH}^{2}$, the seed thus determines the initial sBHAR. In TNG, the seed mass is small enough so that the sBHAR is initially negligible, which, together with a high sSFR at high redshift (e.g. Habouzit et al., 2021, see their figure 4), drives a horizontal trajectory in the $M_{\rm BH}$-$M_{*}$ plane for the galaxy (see the left panel of Fig. 2). The low accretion rate of SMBH also implies that its feedback effect is small, and thus the star formation in the galaxy is regulated only by the stellar feedback. A comparison of the energy release rates via different feedback channels by Weinberger et al. (2018, see their figure 1) indeed shows the domination of stellar feedback in the early time. The first phase is thus referred to as the star-formation dominated phase in this paper.
The top left panel of Fig. 3 shows the specific growth rate of different masses ($M_{\rm h}$, $M_{*}$ and $M_{\rm BH}$) as a function of time. Each rate is averaged over the sample defined in §2. To see the growth of the whole sample during the early phases, the time is measured with respect to $t_{\rm 12}$, the ending point of Phase-1, and scaled by $\tau_{2}+\tau_{3}$, the total time of the following two phases. The exclusion of $\tau_{1}$ from the scaling factor is to avoid the ambiguity caused by the seeding time, and the inclusion of $\tau_{3}$ accounts for the fact that $\tau_{2}$ is too short. During Phase-1, the sSFR follows the sHAR tightly, indicating that the formation of the stellar contents is driven by the gas accretion along with the halo growth and that the $M_{*}$-$M_{\rm h}$ relation has been well established before the seeding of SMBH. The sBHAR is significantly lower compared to the rates of the halo and galaxy, indicating that the early SMBH growth can deviate from the $M_{\rm BH}$-$M_{*}$ relation at $z\approx 0$ in both TNG and observations (e.g. Kormendy & Ho, 2013; Greene et al., 2020; Graham & Sahu, 2023a, b; Habouzit et al., 2021; Zhuang & Ho, 2023). Thus, the growth of SMBH in Phase-1 has to be understood by combining the established environment of the galaxy and the halo, and the specific strategy in the seeding of SMBH.
@@
$\rho_{\rm gas}$ and $c_{\rm s}$ are the gas density and effective sound speed, respectively, within a smoothing kernel around the SMBH, $n_{\rm gas}$ is the number density of gas molecules assuming a mean molecular weight of $\mu=1.2$. Thus, $A_{\rm s}$ depends only on the gas environment around the SMBH and, by construction, is a function of the specific entropy, which is shaped by the dissipation and heating processes. The bottom three panels of Fig. 3 show the evolution of gas mass within different spatial extents. The total mass, and those in hot and cold phases, are shown by curves with different colors. Within $2R_{\rm*,1/2}$, cold gas dominates the total gas mass during Phase-1, suggesting that supernova feedback alone cannot quench the star formation. This is consistent with the cooling rate of non-star-forming cells computed by Weinberger et al. (2018, see their figure 1), which always balances the energy output from stellar feedback in the early time. In the innermost region ($r\lesssim 0.2R_{\rm*,1/2}$) during Phase-1, the cold component also dominates the gas mass, which, together with the high gas density in this region (e.g. Bower et al., 2017), implies that the early growth of SMBH is fed mainly by star-forming gas. The quantitative state of star-forming gas cells, however, depends on the subgrid recipes assumed by TNG, which cannot be estimated analytically. We therefore directly compute the state variables from the simulation data. The gas density, $\rho_{\rm gas}$, is averaged by an SPH kernel (e.g. Weinberger et al., 2017) over a sphere centered on the SMBH and enclosing approximately $256$ cells (defined by the radius BH_Hsml in the particle data of TNG). The effective sound speed, $c_{\rm s}$, which includes the contribution from the Alfvén speed of the magnetic field, is obtained by inserting Eq. (3) using the values of $\rho_{\rm gas}$ estimated as above, and $M_{\rm BH}$ and $\dot{M}_{\rm BH}$ read from the simulation data. The evolutions of $A_{\rm s}$, $\rho_{\rm gas}$ and $c_{\rm s}$ are shown in the top right panel of Fig. 3. An interesting fact is the constant value of $A_{\rm s}$ during Phase-1, which implies that the gas processes around the SMBH are effectively adiabatic. We note that this is an outcome of the specific choice of TNG on the effective equation of state (eEOS) for star-forming cells and the subgrid model for the magnetic field (e.g. Springel & Hernquist, 2003; Vogelsberger et al., 2013; Pillepich et al., 2018b).
@@
Fig. 4 shows the distribution of $\tau_{1}$ directly obtained from the fitting function described in §3, in comparison with the value estimated by Eq. (6). To avoid large fluctuations at a specific snapshot, we use the averaged $A_{\rm s}$ during Phase-1. The contour in the figure shows the distribution of SMBHs and the dot-dashed line shows a power-law fitting. The normalization and power index of the fitting are both close to 1, indicating that the estimation of $\tau_{1}$ by Eq. (6) is accurate. Thus, the transition from star-formation dominated growth to SMBH-accretion dominated growth is a combined consequence of the subgrid model adopted by TNG: the low-mass seeds, the Bondi accretion rate, and the eEOS of star-forming gas.
@@ 4.2 Phase-2: SMBH-accretion dominated growth
The transition point, $t_{\rm 12}$, by definition, is the time when the sBHAR becomes larger than the sSFR. As shown in the top left panel of Fig. 3, the sBHAR and sSFR cross each other at this point. In the $M_{\rm BH}$-$M_{*}$ plane, the trajectory makes a transition at $t_{12}$ from nearly horizontal to nearly vertical, as seen from the example shown in Fig. 2. After this point, the sBHAR continues to increase, reaches a peak, and then declines to a rate comparable to the sSFR. In contrast, the sSFR always declines with time and remains below the sBHAR during Phase-2. The obvious reason for the fast growth of the SMBH in Phase-2 is the Bondi accretion implemented by TNG. For a given gas environment, the BHAR is $\propto M_{\rm BH}^{2}$, so that
@@
where $E_{\rm gas}\equiv E_{\rm in}+E_{\rm k}+E_{\rm v}$ is the total the energy of the gas, defined as the sum of the internal, kinetic, and gravitational potential energies, $E_{\rm cool}$ is the energy loss due to radiative cooling, and $E_{\rm fdbk}$ is the energy input from feedback. The ending of fast SMBH accretion requires a significant reduction in the amount of gas that can fuel the SMBH. To be achieved, the gas of the entire galaxy, in particular that in the inner region, has to be significantly affected by the feedback, because otherwise cold gas clumps that can form in the galaxy will sink into the center and fuel the SMBH. Motived by the flat curve of the total gas mass shown in the bottom center panel of Fig. 3 after $t_{12}$, we express the gas mass at time $t$ around $t_{\rm 23}$ as
@@
where $\eta_{\rm thermal}=0.02$ is the efficiency of the thermal-mode feedback (Weinberger et al., 2018), The radiative cooling is dominated by star-forming gas cells because of their high density and the sensitive dependence of the cooling rate on gas density. Unfortunately, star-forming gas cells are not resolved by the simulation, which prevents us from computing the cooling rate directly. Weinberger et al. (2018) suggested an alternative way to estimate the cooling loss by subtracting the feedback energy deposited to the star-forming gas cells. This gives a value of $\eta_{\rm eff}\approx 10\%$$20\%$ for TNG, as shown in their figure 1. Substituting Eqs. (12) and (4.2) into Eq. (4.2), we can obtain the condition for the ending of the fast SMBH accretion as
@@
The details of the gas ejection by AGN feedback depend on the structure and thermal state of the gas, the coupling of the feedback energy to the gas, and the transfer of energy among gas cells. In TNG, the energy of the thermal-mode feedback is directly deposited to the gas cells surrounding the SMBH and transferred to the outer region via hydrodynamical processes (Weinberger et al., 2017). Thus, the impact of the thermal-mode feedback is inside-out: it significantly reduces the amount of gas surrounding the SMBH, but has smaller effects on outer regions. The three bottom panels of Fig. 6 show the change of the gas mass within different scales. As expected, the gas mass on halo scales, for both the cold and hot components, continues to increase. The mass of the galactic-scale gas stops growing at around $t_{\rm 23}$. Within $0.2R_{\rm*,1/2}$, the cold gas is heated and ejected by the AGN feedback, leading to a reduction of the total gas mass after $t_{\rm 23}$. The temperature of the gas within the galaxy is raised, but the density drops, as shown in the top right panel of Fig. 6. The significant inside-out change of the gas profile during Phase-2 leads to the cease of SMBH accretion, but only a moderate reduction in star formation, as seen from the top left panel of Fig. 6. The evolution path of the galaxy in the $M_{\rm BH}$-$M_{*}$ plane, as shown by the example in Fig. 2, is then pulled back from nearly vertical to a shallower slope. Note that the results obtained here are consistent with the conclusion of Terrazas et al. (2020) that the thermal-mode AGN feedback primarily regulates the growth of the SMBH, but cannot quench the star formation. Instead, the growth of the galaxy enters into a self-regulated phase, as discussed below.

### v3

@@ 3 Evolutionary Phases and Transitions
Figure 2: The phases and transitions in the evolution histories of galaxies. Left panel shows the $_{\rm BH}-MM_{*}$ histories of galaxies. Grey curves represent the simulated histories of 20 individual galaxies randomly selected from our sample (see §2). Dots are the history of an example galaxy, colored blue (red) if it is star-forming (quenched) at a snapshot. Red dashed curve shows the analytical fitting of the history for this example. Four phases are defined according to the fitting function and are labeled at the top of the panel. Three time points, $t_{\rm 12}$, $t_{\rm 23}$ and $t_{\rm 34}\equiv t_{\rm q}$, are used to separate these phases (see §3 for the fitting method and the definition of phases and transitions). Black curve shows the median $_{\rm BH}-MM_{*}$ relation at $z=0$ computed using all galaxies in our sample, with the shaded area indicating the $16^{\rm th}$$84^{\rm th}$ percentiles. Short black line indicates a power law, $M_{\rm BH}\propto M_{*}^{5/3}$, to which the regulation of thermal-mode AGN feedback in Phase-3 tends to drive the galaxy (see §4.3). Right panel shows the cumulative distribution of $\tau_{i}$, the duration of Phase-ii.
The similarity of the trajectories of the quiescent galaxies in Fig. 1 indicates that a universal function can be used to fit the evolution path of each quiescent galaxy in the $_{\rm BH}-MM_{*}$ plane. To quantify the transitions and phases in the growth histories, we fit all snapshots before the quenching time, $t_{\rm q}$, by the following piecewise continuous function:
@@
where $y=\log(M_{\rm BH}/\,{\rm M_{\odot}})$, $x=\log(M_{\rm*}/\,{\rm M_{\odot}})$, and $(x_{0},y_{0},k,a,b)$ are the five free parameters characterizing the evolutionary phases and transitions. The values of the free-parameters are obtained by minimizing the loss function, $\chi^{2}=\sum_{s}(y_{{\rm sim},s}-y_{s})^{2}$, computed using the simulated value ($y_{\rm sim}$) and the value predicted by the fitting function ($y$) at every snapshot, $s$. The optimization is performed using the Levenberg-Marquardt algorithm, which is designed for solving non-linear least squares problems. By design, the fitting function explicitly includes three transition points at which the behavior of the galaxy in the $_{\rm BH}-MM_{*}$ plane changes significantly. The first transition point is $x=x_{0}-b$ at which ${\rm d}y/{\rm d}x=1$. This marks the time when the growth of the SMBH begins to outpace the star formation. The second transition is at $x=x_{0}$, where the rapid growth in the SMBH mass ends. Given that TNG adopts a Bondi accretion upper-bound only by the Eddington limit, the growth of SMBH during $x_{0}-b<x<x_{0}$ can be very fast. This behavior can be captured by a value of $a$ close to zero. For $x\geqslant x_{0}$, the function $y(x)$ is designed to be linear, describing a power-law relation, $M_{\rm BH}\propto M_{*}^{k}$. The final transition is the ending point of the domain of the fitting function, which, by design, is the quenching point of the galaxy. The red dashed curve shows the best-fit function for an example galaxy in the left panel of Fig. 2, with three black hexagons marking the three transition points (see below for the notations), respectively. The fitting captures well all the nonlinear features of the simulated points in this example. Since mergers do not dominate the growth of $M_{\rm BH}$ and $M_{*}$ compared with the accretion of gas and in-situ star formation before quenching (see e.g. figure 7 of Weinberger et al., 2018), the nuances introduced by mergers in the growth of galaxies in the $_{\rm BH}-MM_{*}$ plane do not significantly affect the fitting. For reference, in Appendix A, we quantify the goodness-of-fit for all galaxies in our sample. For most galaxies in our sample, the root-mean-square error is less than $0.1\,{\rm dex}$, and the maximum absolute error is less than $0.3\,{\rm dex}$, indicating that the fittings are reliable for individual galaxies. In addition, we have performed tests by tweaking the fitting function and resampling the data points used in the fitting, and found that the identification of the transition points is robust against such changes.
@@
  • Phase-4: Merger dominated growth. This phase is characterized by the cease of in-situ star formation and SMBH growth. Consequently, the trajectory of a galaxy in the $_{\rm BH}-MM_{*}$ plane is stuck at around a point near the scaling relation at $z=0$, with some occasional “jumps” due to mergers.
@@ 4 Processes Driving the Evolution
Figure 3: Evolutionary history of various properties around the transition point $t_{12}$. Each property describing the history is averaged over our sample (see §2) and shown as a function of time, measured with respect to $t_{12}$ and scaled by the duration $\tau_{2}+\tau_{3}$. The shaded area around each curve shows the $16^{\rm th}$$84^{\rm th}$ percentiles. Each panel shows a set of related properties. Top left panel shows the specific growth rates of of $M_{\rm h}$ (blue), $M_{*}$ (orange) and $M_{\rm BH}$ (green). Top right panel shows the gas properties involved in the definition of Bondi accretion rate (see Eq. 3 and 4.1) in the vicinity of the central SMBH. Three bottom panels show the gas masses defined at different scales: bound to the subhalo (left), within $2R_{\rm*,1/2}$ (middle), and within $0.2R_{\rm*,1/2}$ (right). Total, cold ($\leqslant 10^{5}\,{\rm K}$), and hot ($10^{5}\,{\rm K}$) gas masses are shown by black, blue and red curves, respectively, all scaled by total gas mass at $t_{12}$.
Figure 4: Relation between the estimated value and the simulated value of the Phase-1 duration ($\tau_{1}$). Grey curves, from inner to outer, are contours enclosing 10% – 95% of the sample. Dashed line is a power-law fitting (in linear scale) of the relation. The normalization and power index are both close to 1, indicating that the estimated value well approximates the simulated value. See §4.1 for the details.
@@ 4.1 Phase-1: Star-formation dominated growth
The growth of a galaxy in the first phase is featured by a small mass of the SMBH. In TNG, a seed of SMBH with a mass of $M_{\rm BH}=1.18\times 10^{6}\,{\rm M_{\odot}}$ is placed into a galaxy when the mass of its host halo exceeds $7.38\times 10^{10}\,{\rm M_{\odot}}$ (Weinberger et al., 2018). The subsequent growth of the SMBH is modeled by the Bondi-Hoyle accretion (Weinberger et al., 2017, thereafter Bondi accretion). As the accretion rate of the SMBH modeled in this way scales as $M_{\rm BH}^{2}$, the seed thus determines the initial sBHAR. In TNG, the seed mass is small enough so that the sBHAR is initially negligible, which, together with a high sSFR at high redshift (e.g. Habouzit et al., 2021, see their figure 4), drives a horizontal trajectory in the $_{\rm BH}-MM_{*}$ plane for the galaxy (see the left panel of Fig. 2). The low accretion rate of SMBH also implies that its feedback effect is small, and thus the star formation in the galaxy is regulated only by the stellar feedback. A comparison of the energy release rates via different feedback channels by Weinberger et al. (2018, see their figure 1) indeed shows the domination of stellar feedback in the early time. The first phase is thus referred to as the star-formation dominated phase in this paper.
The top left panel of Fig. 3 shows the specific growth rate of different masses ($_{\rm h}, MM_{*}$ and $M_{\rm BH}$) as a function of time. Each rate is averaged over the sample defined in §2. To see the growth of the whole sample during the early phases, the time is measured with respect to $t_{\rm 12}$, the ending point of Phase-1, and scaled by $\tau_{2}+\tau_{3}$, the total time of the following two phases. The exclusion of $\tau_{1}$ from the scaling factor is to avoid the ambiguity caused by the seeding time, and the inclusion of $\tau_{3}$ accounts for the fact that $\tau_{2}$ is too short. During Phase-1, the sSFR follows the sHAR tightly, indicating that the formation of the stellar contents is driven by the gas accretion along with the halo growth and that the $M_{*}$-$M_{\rm h}$ relation has been well established before the seeding of SMBH. The sBHAR is significantly lower compared to the rates of the halo and galaxy, indicating that the early SMBH growth can deviate from the $_{\rm BH}-MM_{*}$ relation at $z\approx 0$ in both TNG and observations (e.g. Kormendy & Ho, 2013; Greene et al., 2020; Graham & Sahu, 2023a, b; Habouzit et al., 2021; Zhuang & Ho, 2023). Thus, the growth of SMBH in Phase-1 has to be understood by combining the established environment of the galaxy and the halo, and the specific strategy in the seeding of SMBH.
@@
$\rho_{\rm gas}$ and $c_{\rm s}$ are the gas density and effective sound speed, respectively, within a smoothing kernel around the SMBH, $n_{\rm gas}$ is the number density of gas molecules, assuming a mean molecular weight of $\mu=1.2$. Thus, $A_{\rm s}$ depends only on the gas environment around the SMBH and, by construction, is a function of the specific entropy, which is shaped by the dissipation and heating processes. The bottom three panels of Fig. 3 show the evolution of gas mass within different spatial extents. The total mass, and those in hot and cold phases, are shown by curves with different colors. Within $2R_{\rm*,1/2}$, cold gas dominates the total gas mass during Phase-1, suggesting that supernova feedback alone cannot quench the star formation. This is consistent with the cooling rate of non-star-forming cells computed by Weinberger et al. (2018, see their figure 1), which always balances the energy output from stellar feedback in the early time. In the innermost region ($r\lesssim 0.2R_{\rm*,1/2}$) during Phase-1, the cold component also dominates the gas mass, which, together with the high gas density in this region (e.g. Bower et al., 2017), implies that the early growth of SMBH is fed mainly by star-forming gas. The quantitative state of star-forming gas cells, however, depends on the subgrid recipes assumed by TNG, which cannot be estimated analytically. We therefore directly compute the state variables from the simulation data. The gas density, $\rho_{\rm gas}$, is averaged by an SPH kernel (e.g. Weinberger et al., 2017) over a sphere centered on the SMBH and enclosing approximately $256$ cells (defined by the radius BH_Hsml in the particle data of TNG). The effective sound speed, $c_{\rm s}$, which includes the contribution from the Alfvén speed of the magnetic field, is obtained by inserting Eq. (3) using the values of $\rho_{\rm gas}$ estimated as above, and $M_{\rm BH}$ and $\dot{M}_{\rm BH}$ read from the simulation data. The evolutions of $A_{\rm s}$, $\rho_{\rm gas}$ and $c_{\rm s}$ are shown in the top right panel of Fig. 3. An interesting fact is the constant value of $A_{\rm s}$ during Phase-1, which implies that the gas processes around the SMBH are effectively adiabatic. We note that this is an outcome of the specific choice of TNG on the effective equation of state (eEOS) for star-forming cells and the subgrid model for the magnetic field (e.g. Springel & Hernquist, 2003; Vogelsberger et al., 2013; Pillepich et al., 2018b).
@@
Fig. 4 shows the distribution of $\tau_{1}$ directly obtained from the fitting function described in §3, in comparison with the value estimated by Eq. (6). To avoid large fluctuations at a specific snapshot, we use the average $A_{\rm s}$ during Phase-1. The contour in the figure shows the distribution of SMBHs and the dot-dashed line shows a power-law fitting. The normalization and power index of the fitting are both close to 1, indicating that the estimation of $\tau_{1}$ by Eq. (6) is accurate. Thus, from Eq. (6), the transition from star-formation dominated growth to SMBH-accretion dominated growth, with a given definition of the transition, is a combined consequence of the subgrid model adopted by TNG: the low-mass seeds, the Bondi accretion rate, and the eEOS of star-forming gas.
@@ 4.2 Phase-2: SMBH-accretion dominated growth
The transition point, $t_{\rm 12}$, by definition, is the time when the sBHAR becomes larger than the sSFR. As shown in the top left panel of Fig. 3, the sBHAR and sSFR cross each other at this point. In the $_{\rm BH}-MM_{*}$ plane, the trajectory makes a transition at $t_{12}$ from nearly horizontal to nearly vertical, as seen from the example shown in Fig. 2. After this point, the sBHAR continues to increase, reaches a peak, and then declines to a rate comparable to the sSFR. In contrast, the sSFR always declines with time and remains below the sBHAR during Phase-2. The obvious reason for the fast growth of the SMBH in Phase-2 is the Bondi accretion implemented by TNG. For a given gas environment, the BHAR is $\propto M_{\rm BH}^{2}$, so that
@@
where $E_{\rm gas}\equiv E_{\rm in}+E_{\rm k}+E_{\rm v}$ is the total energy of the gas, defined as the sum of the internal, kinetic, and gravitational potential energies, $E_{\rm cool}$ is the energy loss due to radiative cooling, and $E_{\rm fdbk}$ is the energy input from feedback. The ending of fast SMBH accretion requires a significant reduction in the amount of gas that can fuel the SMBH. To be achieved, the gas of the entire galaxy, in particular that in the inner region, has to be significantly affected by the feedback, because otherwise cold gas clumps that can form in the galaxy will sink into the center and fuel the SMBH. Motivated by the flat curve of the total gas mass shown in the bottom center panel of Fig. 3 after $t_{12}$, we express the gas mass at time $t$ around $t_{\rm 23}$ as
@@
where $\eta_{\rm thermal}=0.02$ is the efficiency of the thermal-mode feedback (Weinberger et al., 2018). To exclude the contribution from mergers to $\dot{M}_{\rm BH}$ in the estimation of $E_{\rm agn}$, we take the field BH_CumEgyInjection_QM from the catalog of TNG to obtain the feedback energy injected by the BH in the thermal mode. The radiative cooling is dominated by star-forming gas cells because of their high density and the sensitive dependence of the cooling rate on gas density. Unfortunately, star-forming gas cells are not resolved by the simulation, which prevents us from computing the cooling rate directly. Weinberger et al. (2018) suggested an alternative way to estimate the cooling loss by subtracting the feedback energy deposited to the star-forming gas cells. This gives a value of $\eta_{\rm eff}\approx 10\%$$20\%$ for TNG, as shown in their figure 1. Substituting Eqs. (12) and (4.2) into Eq. (4.2), we can obtain the condition for the ending of the fast SMBH accretion as
@@
The details of the gas ejection by AGN feedback depend on the structure and thermal state of the gas, the coupling of the feedback energy to the gas, and the transfer of energy among gas cells. In TNG, the energy of the thermal-mode feedback is directly deposited to the gas cells surrounding the SMBH and transferred to the outer region via hydrodynamical processes (Weinberger et al., 2017). Thus, the impact of the thermal-mode feedback is inside-out: it significantly reduces the amount of gas surrounding the SMBH, but has smaller effects on outer regions. The three bottom panels of Fig. 6 show the change of the gas mass within different scales. As expected, the gas mass on halo scales, for both the cold and hot components, continues to increase. The mass of the galactic-scale gas stops growing at around $t_{\rm 23}$. Within $0.2R_{\rm*,1/2}$, the cold gas is heated and ejected by the AGN feedback, leading to a reduction of the total gas mass after $t_{\rm 23}$. The temperature of the gas within the galaxy is raised, but the density drops, as shown in the top right panel of Fig. 6. The significant inside-out change of the gas profile during Phase-2 leads to the cease of SMBH accretion, but only a moderate reduction in star formation, as seen from the top left panel of Fig. 6. The evolutionary path of the galaxy in the $_{\rm BH}-MM_{*}$ plane, as shown by the example in Fig. 2, is then pulled back from nearly vertical to a shallower slope. Note that the results obtained here are consistent with the conclusion of Terrazas et al. (2020) that the thermal-mode AGN feedback primarily regulates the growth of the SMBH, but cannot quench the star formation. Instead, the growth of the galaxy enters into a self-regulated phase, as discussed below.